At a furniture store, a chair was sold for Php 475 less than a side table and Php 278 more than a kitchen dish cabinet. The total cost of three items qas Php 4175. What is the selling price of the side table?
Answer by richwmiller(17219)

c=t-475
t=c+475
c=k+278
k=c-278
t=c+475
c+t+k=4175
substitute